# Harnessing Multiple Large Language Models: A Survey on LLM Ensemble

### v4

@@ Abstract
Behind this research enthusiasm, however, we can identify two main aspects: 1) The performance concerns: The direct out-of-the-box capability of LLMs (from zero-shot inference) and their indirect out-of-the-box capability (from in-context-learning few-shot inference) still raise performance worries, including accuracy, hallucinations, and misalignment with human intent, among others; 2) The varying strengths and weaknesses of LLMs, each with different inference costs: Due to differences in architecture, parameter size, tokenization, dictionary, training data, and methodology, these LLMs exhibit substantial variability and their responses can differ significantly. With the above two aspects in mind and drawing on the spirit of Ensemble Learning, it is natural to consider that, for each task query, rather than persistently relying on a single LLM based on public rankings or other criteria, it might be more advantageous to simultaneously consider multiple LLM candidates (usable out-of-the-box) and harness their distinct strengths. This is exactly what the recently emerging field of LLM Ensemble explores.
@@ 2.2 Related Problems
LLM Merging, also known as LLM Fusion Yang et al. (2024), integrates parameters from various LLMs to construct a universal model without requiring original training data or extensive computation. It is highly relevant to LLM Ensemble, as both promote knowledge fusion and transfer.
LLM Collaboration leverages the distinct strengths of each model to approach tasks with increased flexibility Du et al. (2023); Lu et al. (2024a). Unlike LLM Ensemble where models are employed with equal status to directly faced to user queries, the collaboration approach assigns distinct roles to each LLM, exchanging response information to enhance task resolution.
Weak Supervision Zhang et al. (2021); Chen et al. (2023c), also known as Learning from Crowds Chen et al. (2021, 2022), uses weak labels to perform information aggregation Chen et al. (2023b) (corresponding non-cascade ensemble after inference in LLM Ensemble) or directly train a classifier Zhang et al. (2021), yet most methods focus on classification, rather than general generation tasks.
@@ 3 Methodology
for tree=
grow=east,
reversed=true,
anchor=base west,
parent anchor=east,
child anchor=west,
base=left,
font=,
rectangle,
draw,
rounded corners,align=left,
minimum width=2.5em,
inner xsep=4pt,
inner ysep=1pt,
where level=1text width=5em,fill=blue!10,
where level=2text width=5em,font=,fill=pink!30,
where level=3font=,yshift=0.26pt,fill=yellow!20,
[LLM
Ensemble
methods,fill=green!20
[(a) Ensemble-before-
inferece (Sec. 3.1),text width=7.6em
[(a1) Pretrained router,text width=7.8em
[Classification-based: Shnitzer et al. (2023) / Srivatsa et al. (2024) / Ong et al. (2024) / Hybrid-LLM Ding et al. (2024) /
SelectLLM Maurya et al. (2024) / Routoo Mohammadshahi et al. (2024)
[Reward-based: Zooter Lu et al. (2024b) / Bench-CoE Wang et al. (2024) / Li (2025) / MetaLLM Nguyen et al. (2024) /
TO-Router Stripelis et al. (2024)
[Assignment-based: FORC Šakota et al. (2024) / HomoRouter Mu et al. (2024)
[(a2) Non-pretrained router,text width=9.8em [Eagle Zhao et al. (2024) / PickLLM Sikeridis et al. (2024) / Blending Lu et al. (2024c)
[(b) Ensemble-during-
inference (Sec. 3.2),text width=7.6em
[(b1) Token-level ensemble,text width=9.6em
[Aggregation-based: 1) Vocabulary alignment: GaC Yu et al. (2024) / DeePEn Huang et al. (2024) / EVA Xu et al. (2024)
/ 2) Others: UniTe Yao et al. (2024) / PackLLM Mavromatis et al. (2024) / CDS Jin et al. (2024) / Li et al. (2024b)
[Selection-based: CITER Zheng et al. (2025)
[(b2) Span-level ensemble, text width=9.6em
[Cool-Fusion Liu et al. (2024b) / SweetSpan Xu et al. (2025) / SpecFuse Lv et al. (2024b)
[(b3) Process-level ensemble, text width=10.1em
[LE-MCTS Park et al. (2024)
[(c) Ensemble-after-
inference (Sec. 3.3),text width=7.6em
[(c1) Non-cascade,text width=6.4em
[Selection-based: Agent-Forest Li et al. (2024a) / Smoothie Guha et al. (2024) / MoRE Si et al. (2023)]
[Selection-then-regeneration: LLM-Blender Jiang et al. (2023) / URG Lv et al. (2024a) / LLM-TOPLA Tekin et al. (2024) ]
[(c2) Cascade, text width=5.4em
[Unsupervised: EcoAssistant Zhang et al. (2023) / Yue et al. (2024) / Model Cascading Varshney and Baral (2022) /
neural caching Ramírez et al. (2023) / Cascade Routing Dekoninck et al. (2024) ]
[Supervised: Jitkrittum et al. (2024) / Gupta et al. (2024) / FrugalGPT Chen et al. (2023a) / AutoMix Aggarwal et al. (2023) /
DER Hu et al. (2024a) / ]
@@ 3.1 Ensemble Before Inference
FORC Šakota et al. (2024) provides a variety of assignment strategies, including both cost-insensitive strategies and cost-sensitive strategies. Users can choose the appropriate strategy according to different needs and constraints to achieve the best balance between cost and performance. In the retrieval-augmented generation (RAG) scenario, given a set of homogeneous search tools and user queries, HomoRouter Mu et al. (2024) is trained to predict the score of the LLM calling each tool to process the query, and then the query is assigned to the best tool according to different allocation strategies.

### v5

@@ 1 Introduction
Behind this research enthusiasm, however, we can identify two main aspects: 1) The performance concerns: The direct out-of-the-box capability of LLMs (from zero-shot inference) and their indirect out-of-the-box capability (from in-context-learning few-shot inference) still raise performance worries, including accuracy, hallucinations, and misalignment with human intent, among others; 2) The varying strengths and weaknesses of LLMs, each with different inference costs: Due to differences in architecture, parameter size, tokenization, dictionary, training data, and methodology, these LLMs exhibit substantial variability and their responses can differ significantly. With the above two aspects in mind and drawing on the spirit of Ensemble Learning Mohammed and Kora (2023); Zhou (2021); Dong et al. (2020), it is natural to consider that, for each task query, rather than persistently relying on a single LLM based on public rankings or other criteria, it might be more advantageous to simultaneously consider multiple LLM candidates (usable out-of-the-box) and harness their distinct strengths. This is exactly what the recently emerging field of LLM Ensemble explores.
@@ 2.2 Related Problems
LLM Merging, also known as LLM Fusion Yang et al. (2024, 2025); Ruan et al. (2025), integrates parameters from various LLMs to construct a universal model without requiring original training data or extensive computation. It is highly relevant to LLM Ensemble, as both promote knowledge fusion and transfer.
LLM Collaboration leverages the distinct strengths of each model to approach tasks with increased flexibility Du et al. (2023); Lu et al. (2024a); Chen et al. (2025); Tran et al. (2025); Zhang et al. (2025). Unlike LLM Ensemble where models are employed with equal status to directly faced to user queries, the collaboration approach assigns distinct roles to each LLM, exchanging response information to enhance task resolution.
Weak Supervision Zhang et al. (2021); Chen et al. (2023c); Zhang et al. (2022); Ratner et al. (2017), sometimes referred to as Learning from Crowds Chen et al. (2021, 2022); Zhang (2022), uses weak labels contributed from multiple weak annotation sources to perform information aggregation Chen et al. (2023b); Zheng et al. (2017) (corresponding non-cascade ensemble after inference in LLM Ensemble) or directly train a classifier Zhang et al. (2021), yet most methods focus on classification rather than general open-ended generation tasks.
@@ 3.1 Ensemble Before Inference
FORC Šakota et al. (2024) provides a variety of assignment strategies, including both cost-insensitive strategies and cost-sensitive strategies. Users can choose the appropriate strategy according to different needs and constraints to achieve the best balance between cost and performance. In the retrieval-augmented generation (RAG) scenario, given a set of homogeneous search tools and user queries, HomoRouter Mu et al. (2024) is trained to predict the score of the LLM calling each tool to process the